OpenStudy (anonymous):

Please help me solve this simultaneous equation 3x-2y=10 2x+5y=15

OpenStudy (anonymous):

This is a system of equations, right?

OpenStudy (anonymous):

and does it want you to solve by elimination, substitution, etc?

OpenStudy (anonymous):

@Jneil ?

OpenStudy (anonymous):

Am not sure i guess.

OpenStudy (anonymous):

Okay, well I used the elimination method and this is what I got.. \[x= \frac{ 80 }{ 19 }, y=\frac{ 25 }{ 19 }\]

OpenStudy (anonymous):

Is this a multiple choice question?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Do you have to show work?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Okay, give me a minute!

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Multiply each equation by the value that makes the coefficients of y equal. This value is found by dividing the least common multiple of the coefficients of y by the current coefficient. In this case, the least common multiple is 10. 5⋅(3x−2y=10) 2⋅(2x+5y=15) Multiply each equation by the value that makes the coefficients of y equal. This value is found by dividing the least common multiple of the coefficients of y by the current coefficient. In this case, the least common multiple is 10. 5⋅(3x−2y)=5(10) 2⋅(2x+5y)=2(15) Multiply 5 by each term inside the parentheses. 5⋅(3x−2y)=50 2⋅(2x+5y)=2(15) Multiply 5 by each term inside the parentheses. 15x−10y=50 2⋅(2x+5y)=2(15) Multiply 2 by each term inside the parentheses. 15x−10y=50 2⋅(2x+5y)=30 Multiply 2 by each term inside the parentheses. 15x−10y=50 4x+10y=30

OpenStudy (anonymous):

Add the two equations together to eliminate y from the system.

OpenStudy (anonymous):

Am not getting it

OpenStudy (anonymous):

Divide each term in the equation by 19. x=80/19 Substitute the value found for x into the original equation to solve for y. 15(80/19)−10y=50 Multiply 15 by each term inside the parentheses. 1200/19−10y=50 Move all terms not containing y to the right-hand side of the equation. −10y=−250/19 Divide each term in the equation by −10. y=25/19 This is the final solution to the independent system of equations. Answer: x= 80/19, y=25/19 Remember, I used the elimination method.

OpenStudy (anonymous):

Ok am gonna try it

OpenStudy (anonymous):

You add the two together and are trying to get x along and then y alone, using the simplest equation.

OpenStudy (anonymous):

Can you use the substitution method please.